# Feynman’s argument shows that GWs would not transfer energy to dust or a perfect fluid because they are frictionless, but there would be an energy transfer to a viscous fluid. In 1966, Hawking [17] considered GW propagation in different cosmological models. It follows that

### v1

@@ 1 Introduction
This essay is structured as follows: Section 2 introduces matter shell theory, GW damping, and GW heating. Section 3 delves into the astrophysical contexts where GW heating and GW damping is significant. Finally, the concluding section summarizes the key points.
@@ 2 Matter shell around the GW source
2.1 Bondi-Sachs formalism
2.1.1 Null coordinates system
To naturally determine the GW from different astrophysical sources we foliate the space-time by outgoing null hypersurfaces. These null hypersurfaces emanate from the timelike world tube (3-surface) $\Gamma$ which we construct by transporting a convex 2-surface along a timelike vector field. We uniquely describe these 2-surface by choosing the angular coordinates ($x^{A},A=2,3$, and $x^{2}=\theta,x^{3}=\phi$). We also describe $\Gamma$ by the coordinates $(u,x^{A})$ where $u$ is the time-coordinate along $\Gamma$. Furthermore, we label the hypersufaces by the coordinates $x^{\alpha}=(u,r,x^{A})$, $\alpha=0,1,2,3$ where $x^{0}=u=const.$ and $x^{2}=r$ is the surface area coordinate at future null infinity $\mathcal{I^{+}}$.
2.1.2 Bondi-Sachs metric
The Bondi-Sachs metric in null coordinates takes the following form
$ds^{2}=$displaystyle-- $\left[e^{2\beta}\left(1+\frac{W}{r}\right)-r^{2}h_{AB}U^{A}U^{B}\right]du^{2}-2e^{2\beta}dudr$ (3)
@@ 2.1.2 Bondi-Sachs metric
where $h^{AB}h_{BC}=\delta^{A}_{B}$ and $det(h_{AB})=det(q_{AB})$, with $q_{AB}$ being a unit sphere metric and $h_{AB}$ the conformal (angular) metric, The spin-weighted field is given by $U=U^{A}q_{A}$, with $q_{A}$ being the complex dyad. The expansion of the null cone is measured by $e^{2\beta}$ with $\beta$ (through out, this $\beta$ is not the same as that used for the indexes) representing the expansion of the light rays propagating outwards from $\Gamma$ to $\mathcal{I^{+}}$ For a Schwarzschild spacetime, $W=-2M$. We define the complex quantity $J$ by
$J=q^{A}q^{B}h_{AB}/2.\$ (4)
For the Schwarzschild spacetime, we have $J$ and $U$ being zero and thus they can be regarded as a measure of the deviation from spherical symmetry, and in addition, they contain all the dynamic content of the gravitational field in the linearized regime [5]. Usually we can describe this spacetime by $\beta=0$ and $W=-2M$, or by $\beta=\beta_{c}(\mbox{constant})$ and $W=(e^{2\beta_{c}}-1)r-2M$.
2.1.3 Field equations for the Bondi-Sachs metric
In Bondi-Sachs formalism the Einstein field equations
$R_{\alpha\beta}=8\pi(T_{\alpha\beta}-\frac{1}{2}g_{\alpha\beta}T),$ (5)
are categories as [5]:
  • the hypersurface equations $R_{rr}$, $q^{A}R_{rA}$, $h^{AB}R_{AB}$ for $\beta$, $U$, and $W$ respectively,
  • the evolution equations $q^{A}q^{B}R_{AB}$ for $J$, and
  • the constraints equations $R_{0\beta}$.
The Ricci tensor are used to derive these equations. The evolution equations really describe the propagation of GW throughout spacetime. The key about the constraints equations is that if they are satisfied at the inner boundary of $\Gamma$, then they will be satisfied throughout the entire spacetime.
@@ 2.2 Thin matter shell problem
We consider a matter distribution in the form of a thin, low-density shell surrounded by a GW source as shown in Figure 1. The shell is spherical at $r=const.$ and we allow the density $\rho$ to vary around the shell to make the problem generally not spherically symmetric.
Figure 1: A pictorial representation of a GW source surrounded by a spherical shell of mass $M_{S}$ located between $r=r_{0}$ and $r=r_{0}+\Delta$, (r is the distance from the source).
A solution to the Einstein equations can be constructed using linearized perturbation theory, see [5]. Furthermore, it is found that the dust shell causes the gravitational wave to be modified both in magnitude and phase, but without any energy being transferred to or from the dust [5].
The introduction of a spherical shell around the GW source of mass $M_{S}$, radius $r_{0}$ and thickness $\Delta$ modifies the wave strain to:
$H=-2\nu^{2}\sqrt{6}C_{40}\left(1+\frac{2M_{S}}{r_{0}}+\frac{2iM_{S}}{r_{0}^{2}\nu}+\frac{iM_{S}e^{-2ir_{0}\nu}}{2r_{0}^{2}\nu}+{\mathcal{O}}\left(\frac{M_{S}\Delta}{r^{2}_{0}},\frac{M_{S}}{r_{0}^{3}\nu^{2}}\right)\right),$ (6)
where $\nu=2\pi f$ and $C_{40}$ is a constant. Notice that each of the terms containing $M_{S}$ in Eq. 6 represents a correction to the wave strain in the absence of the shell. Furthermore, we have the following observations:
@@
  • The presence of $e^{4\pi ir_{0}f}$ in the third correction term describes a change in the magnitude of $H$ and is due to an incoming wave modifying the geometry near the source and thus the inspiral rate.
@@ 2.3.1 GW damping and GW heating
To see the damping of the GWs inside the viscous shell, let $H$ denote the magnitude of the GWs rescaled to account for a $1/r$ falloff, and let $r_{o}$ and $r_{i}$ represent the outer and inner radii of the shell, respectively. As $r_{o}$ tends to infinity, the magnitude of the GWs can be expressed as
@@ 3.1.1 CCSNe as a dust shell
For a dust shell surrounding a GW emission, modifications to the GW occur in both magnitude and phase [6]. In [26], this analysis was extended to the GW event associated with a CCSNe, treated the protoneutron star (PNS) as a dust shell surrounding a GW emission and, again, it was shown that modifications to the GW occur in both magnitude and phase. The results were shown to be potentially astrophysically relevant with the modification measurable for certain parameters. The magnitude was shown to be proportional to $\lambda/r_{i}$, with $\lambda$ the GW wavelength and $r_{i}$ the inner radius of the matter shell. It was also shown that a burst of GWs can, in principle, lead to echoes [26], although the time difference between the echo and the actual wave and the relative amplitude of the echo to the main signal would mean the echo would be buried in main signal and hence not discernible.
@@ 3.1.2 Viscosity in the core collapse environment
The paper then looked at the astrophysical application of CCSNe. A number of CCSNe models with different parameters were proposed, and significant viscous damping of GWs was predicted for many cases. GW generation in CCSNe involves a number of different physical processes, each with GW output at a different frequency. It may be that a GW observation of a CCSNe event would see only the higher frequencies, with the lower ones completely damped out.
@@ 3.1.3 GW Heating in CCSNe
Following on the work in [21] we then consider whether there is any heating effect on the PNS owing to the viscous damping (in preparation). For realistic scenarios it seems that GW heating of the PNS is a likely scenario. One of the consequences of such heating is that, in the case of a shock wave stalling in a post-bounce phase, GW heating may produce enough energy to resuscitate the shock wave, hence ensuring the PNS results in a CCSNe where it otherwise may not have.
@@ 3.2.1 Cosmological scenario
An illustration follows by considering a hypothetical scenario where GWs propagate from a single source subsequent to the epoch of cosmic inflation in a radiation dominated Einstein-de Sitter (EdS) universe with $a(t)\propto t^{1/2}$. At the end of inflation at $t=10^{-32}$s and $T=10^{27}$ to $10^{28}$. According to [23] (Section 28.1), $T$ is proportional to $1/a(t)$. Thus, at the end of inflation, we take $a=10^{-27}$. The inverse of the Hubble rate does not change during inflation and is taken as $1/\mathcal{H}(=a/\dot{a})=10^{-35}$s, which leads to the horizon scale as $c/\mathcal{H}=3\times 10^{-27}$m.
@@
A estimate for viscosity, from Weinberg 1971 [32], using Eqs (3.20–3.22), rewritten with $c\neq 1$, we get
@@
If we consider a thin shell scenario where $r_{i},r_{o}$ are much larger than the wavelength $\lambda$, solving Eq. (2) for $r\nu\gg 1$ and rewriting in terms of $t$, we have
@@
in geometric units. Now, $\eta$ behaves as $T^{2}$, i.e. as $1/a^{2}$ and further, we are in the radiation dominated era with $a(t)\propto t^{1/2}$. Thus $\eta$ behaves as $1/t$ and we have $dH/dt=-8\pi\eta_{i}cHt_{i}/t$ with $\eta_{i}=9.09\times 10^{22}$ and $t_{i}=10^{-32}$. Let $A=8\pi\eta_{i}ct_{i}=6.9$, then integrating we get
@@
This example shows that if $\eta$ is as large as the value in Eq. (14) then the damping effect is so rapid that it may occur before the quark epoch is reached. Repeating the calculation for the upper estimate of the viscosity of a quark-gluon plasma, i.e., $\eta\leq 5\times 10^{11}$ kg/m/s (see [29]), further damping during the quark-gluon epoch appears to be inconsequential.
Alternatively, considering that the wavelength is in the range of $3\times 10^{-22}$ to $3\times 10^{-10}\,$m while the horizon scale is $c/H=3\times 10^{-27}$ m, the case for $r_{i},r_{o}\ll\lambda$ can be justified and Eq. (9) becomes applicable. Using the same formulation, significant damping can then occur for even small values of $\eta_{i}$. As opposed to the first scenario where the value of $\eta$ dominates damping, this effect is present in geometries where the $(\lambda/r_{i})$ ratio is large, which amplifies the value of the negative exponent in Eq. (9).
As an example, suppose that $r_{i}=3\times 10^{-27}$ m (a pGW source the size of the horizon scale) with $\lambda=3\times 10^{-22}$ m (the lower wavelength limit), then $H_{o}=0.5H_{i}$ for $\eta=2.0\times 10^{25}$ kg/m/s. This value is much smaller than that of Eq. (14), and is of the same order as values considered for CCSNe. The timescale of the damping is $10^{-35}$ s, which is much shorter than the timescale of inflation.
@@ 4 Conclusion
This essay delves into the astrophysical contexts where GW damping and GW heating are highly relevant. Specifically, we examine the effects of a matter shell enveloping the GW source. Initially, we analyze a low-density dust shell, providing the modified GW strain equation (Eq. 6). Expanding our discourse, we explore the implications of a viscous shell encircling the GW source. Energy dissipation occurs within this shell as GWs traverse it, leading to GW damping and heating of the shell.

### v2

@@ 1 Introduction
This paper is structured as follows: Section 2 introduces matter shell theory, GW damping, and GW heating. Section 3 delves into the astrophysical contexts where GW heating and GW damping is significant. Finally, the concluding section summarizes the key points.
@@ 2 Matter shell around the GW source
To naturally determine the GW from different astrophysical sources we foliate the space-time by outgoing null hypersurfaces. These null hypersurfaces emanate from the timelike world tube (3-surface) $\Gamma$ which we construct by transporting a convex 2-surface along a timelike vector field. We uniquely describe the 2-surface by choosing angular coordinates ($x^{A},A=2,3$, e.g. in spherical polars $x^{2}=\theta,x^{3}=\phi$). We also describe $\Gamma$ by the coordinates $(u,x^{A})$ where $u$ is the time-coordinate along $\Gamma$. Furthermore, we label the null hypersufaces by the coordinates $x^{\alpha}=(u,r,x^{A})$, $\alpha=0,1,2,3$ where $x^{0}=u=const.$ and $x^{1}=r$ is the surface area coordinate.
Then the Bondi-Sachs metric in null coordinates takes the following form
$ds^{2}=$displaystyle-- $\left[e^{2\beta}\left(1+\frac{W_{c}}{r}\right)-r^{2}h_{AB}U^{A}U^{B}\right]du^{2}-2e^{2\beta}dudr$ (3)
@@
where $h^{AB}h_{BC}=\delta^{A}_{B}$ and $\det(h_{AB})=\det(q_{AB})$, with $q_{AB}$ being a unit sphere metric (e.g. $q_{AB}dx^{A}dx^{B}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$) and $h_{AB}$ the conformal (angular) metric. We introduce spin-weighted fields $U=U^{A}q_{A}$ and $J=q^{A}q^{B}h_{AB}/2$, where $q^{A}$ is a complex dyad, e.g. $q^{A}=(1,i/\sin\theta)$; see more details in [13, 14, 11]. The use of $J$ and $U$ leads to significant simplifications of the Einstein equations, and also enables the use of a simple separation of variables ansatz in Eqn. (4) below. Note that for the Schwarzschild spacetime, we have $J$ and $U$ being zero, and thus $J,U$ can be regarded as a measure of the deviation from spherical symmetry.
In the following sections, we model the matter shell surrounding the GW source. Section 2.1 examines a thin, low-density shell, while Section 2.2 focuses on the viscous shell around the GW source.
@@ 2.1 Thin matter shell problem
We consider a static matter distribution as a thin, low-density shell surrounding a GW source, illustrated in Figure 1. The shell is spherical at $r=const.$, with the density $\rho$ allowed to vary around the shell, making the problem generally not spherically symmetric.
Figure 1: A pictorial representation of a GW source surrounded by a spherical shell of mass $M_{S}$ located between $r=r_{0}$ and $r=r_{0}+\Delta$ (where $r$ is the distance from the source).
We make the ansatz of small quadrupolar perturbations about Minkowski spacetime with the metric quantities $(\beta,U,W,J)$ taking the form
$\beta=$ $\Re(\beta^{[2,2]}(r)e^{i\nu u}){}_{0}Z_{2,2}\,,\;\;U=\Re(U^{[2,2]}(r)e^{i\nu u}){}_{1}Z_{2,2}\,,$
$W_{c}=$ $\Re(W_{c}^{[2,2]}(r)e^{i\nu u}){}_{0}Z_{2,2}\,,\;\;J=\Re(J^{[2,2]}(r)e^{i\nu u}){}_{2}Z_{2,2}\,,$ (4)
with the superfix [2,2] indicating a coefficient of ${}_{s}Z_{2,2}$. Note that the perturbations oscillate in time with frequency $\nu/(2\pi)$. The quantities ${}_{s}Z_{\ell,m}$ are spin-weighted spherical harmonic basis functions related to the usual ${}_{s}Y_{\ell,m}$ as specified in [12, 11]. Notably, ${}_{0}Z_{\ell,m}$ are real, allowing the description of the metric quantities $\beta$ and $W_{c}$ (which are real) without mode-mixing. However, for $s\neq 0$ ${}_{s}Z_{2,2}$ is generally complex. A general solution may be constructed by summing over the $(\ell,m)$ modes.
As shown in previous work [12, 5], solving the vacuum Einstein equations under the condition of no incoming radiation leads to
$\beta^{[2,2]}={}$ $b_{0},$
$W_{c}^{[2,2]}={}$ $4i\nu b_{0}-2\nu^{4}C_{40}-2\nu^{2}C_{30}+\frac{4i\nu C_{30}-2b_{0}}{r}+\frac{4i\nu^{3}C_{40}}{r}+\frac{12\nu^{2}C_{40}}{r^{2}}$
$-\frac{12i\nu C_{40}}{r^{3}}-\frac{6C_{40}}{r^{4}},$
$U^{[2,2]}={}$ $\frac{\sqrt{6}(-2i\nu b_{0}+\nu^{4}C_{40}+\nu^{2}C_{30})}{3}+\frac{2\sqrt{6}b_{0}}{r}+\frac{2\sqrt{6}C_{30}}{r^{2}}-\frac{4i\nu\sqrt{6}C_{40}}{r^{4}}$
$-\frac{3\sqrt{6}C_{40}}{r^{4}},$
$J^{[2,2]}={}$ $\frac{2\sqrt{6}(2b_{0}+i\nu^{3}C_{40}+i\nu C_{30})}{3}+\frac{2\sqrt{6}C_{30}}{r}+\frac{2\sqrt{6}C_{40}}{r^{3}},$ (5)
[…]
${H}_{M0}=\Re(\mathbf{H}_{M0}\exp(i\nu u))~{}{}_{2}Z_{2,2}~{}~{}\textrm{with}~{}~{}\mathbf{H}_{M0}=-2\sqrt{6}\nu^{2}C_{40}.$ (6)
It is found that the dust shell modifies the gravitational wave both in magnitude and phase, without any energy being transferred to or from the dust [5]. The introduction of a spherical shell around the GW source of mass $M_{S}$, radius $r_{0}$ and thickness $\Delta$ modifies the wave strain to:
${H}=\Re\left(\mathbf{H}_{M0}\left(1+\frac{2M_{S}}{r_{0}}+\frac{2iM_{S}}{r_{0}^{2}\nu}+\frac{iM_{S}e^{-2ir_{0}\nu}}{2r_{0}^{2}\nu}+{\mathcal{O}}\left(\frac{M_{S}\Delta}{r^{2}_{0}},\frac{M_{S}}{r_{0}^{3}\nu^{2}}\right)\right)\exp(i\nu u)\right)~{}{}_{2}Z_{2,2}.$ (7)
Notice that in the absence of the shell ($M_{S}=0$), $H=H_{M0}$ so that the GW is unmodified. Each term containing $M_{S}$ in Eqn. (7) represents a correction to the GW due to the shell:
@@
  • The presence of $e^{-4\pi ir_{0}f}$ in the third correction term describes a change in the magnitude of $H$ and is due to an incoming wave modifying the geometry near the source and thus the inspiral rate.
@@ 2.2.1 GW damping and GW heating
To see the damping of the GWs inside the viscous shell, let $H$ denote the magnitude of the GWs rescaled to account for a $1/r$ falloff, and let $r_{o}$ and $r_{i}$ represent the outer and inner radii of the shell, respectively. As $r_{o}$ tends to infinity, the magnitude of the GWs can be expressed as [7]:
@@ 3.1.1 CCSNe as a dust shell
For a dust shell surrounding a GW emission, modifications to the GW occur in both magnitude and phase [26]. In [6], this analysis was extended to the GW event associated with a CCSNe, treated the PNS as a dust shell surrounding a GW emission and, again, it was shown that modifications to the GW occur in both magnitude and phase. The results were shown to be potentially astrophysically relevant with the modification measurable for certain parameters. The magnitude was shown to be proportional to $\lambda/r_{i}$, with $\lambda$ the GW wavelength and $r_{i}$ the inner radius of the matter shell. It was also shown that a burst of GWs can, in principle, lead to echoes [6], although the time difference between the echo and the actual wave and the relative amplitude of the echo to the main signal would mean the echo would be buried in the main signal and hence not discernible.
@@ 3.1.2 Viscosity in the core collapse environment
The paper then looked at the astrophysical application of CCSNe. A number of CCSNe models with different parameters were proposed. Detectable frequencies of GWs expected from CCSNe, range from 100 to 1300 Hz. Estimates for the radius of the core of the PNS range from 10 to 30 km. The GW energy released from a CCSNe is expected to be approximately $10^{43}-10^{46}$ erg, with the density at core bounce expected to be approximately $3.2\times 10^{14}$ kg/m3 and the dynamic viscosity ranges from $10^{23}$ to $10^{25}$ kg/m/s. Significant viscous damping of GWs was predicted for many cases ranging from a discernible 10% to complete damping. GW generation in CCSNe involves a number of different physical processes, each with GW output at a different frequency. It may be that a GW observation of a CCSNe event would see only the higher frequencies, with the lower ones completely damped out.
@@ 3.1.3 GW Heating in CCSNe
Following on the work in [8] we then consider whether there is any heating effect on the PNS owing to the viscous damping (in preparation). We find
$38\,\mbox{K}\leq(T-T_{0})\leq 7.0\times 10^{20}\,\mbox{K}\,,$ (14)
so that, depending on the actual values of physical parameters, the significance of the effect varies from very high to none. For realistic scenarios, it seems that GW heating of the PNS is a likely scenario. One of the consequences of such heating is that, in the case of a shock wave stalling in a post-bounce phase, GW heating may produce enough energy to resuscitate the shock wave, hence ensuring the PNS results in a CCSNe where it otherwise may not have.
@@ 3.2.1 Cosmological scenario
An illustration follows by considering a hypothetical scenario where GWs propagate from a single source subsequent to the epoch of cosmic inflation in a radiation-dominated Einstein-de Sitter (EdS) universe with $a(t)\propto t^{1/2}$. At the end of inflation at $t=10^{-32}$s and $T=10^{27}$ to $10^{28}$. According to [31, Section 28.1], $T$ is proportional to $1/a(t)$. Thus, at the end of inflation, we take $a=10^{-27}$. Using the start of inflation as $10^{-36}$s with the Hubble rate not changing, due to exponential expansion, the inverse Hubble rate is taken as $1/\mathcal{H}(=a/\dot{a})=10^{-36}$s, which leads to the horizon scale as $c/\mathcal{H}=3\times 10^{-28}$m.
@@
An estimate for viscosity, from Weinberg 1971 [29], using Eqns (3.20–3.22), rewritten with $c\neq 1$, we get  [^1] Even though there are more recent results of $\eta$, we take the value derived from Weinberg as a baseline for an extreme viscosity under ultra-high energy conditions to illustrate the damping effect.
@@
If we consider a thin shell scenario where $r_{i},r_{o}$ are much larger than the wavelength $\lambda$, solving Eqn. (2) for $r\nu\gg 1$ and rewriting in terms of $t$, we have
@@
in geometric units. Now, $\eta$ behaves as $T^{2}$, i.e. as $1/a^{2}$ and further, we are in the radiation-dominated era with $a(t)\propto t^{1/2}$. Thus $\eta$ behaves as $1/t$ and we have $dH/dt=-8\pi\eta_{i}cHt_{i}/t$ with $\eta_{i}=9.09\times 10^{22}$ and $t_{i}=10^{-32}$. Let $A=8\pi\eta_{i}ct_{i}=6.9$, then integrating we get
@@
This example shows that if $\eta$ is as large as the value in Eqn. (16) then the damping effect is so rapid that it may occur before the quark epoch is reached. Repeating the calculation for the upper estimate of the viscosity of a quark-gluon plasma, i.e., $\eta\leq 5\times 10^{11}$ kg/m/s (see [32]), further damping during the quark-gluon epoch appears to be inconsequential.
Alternatively, considering that the wavelength is in the range of $3\times 10^{-22}$ to $3\times 10^{-10}\,$m while the horizon scale is $c/\mathcal{H}=3\times 10^{-27}$ m, the case for $r_{i},r_{o}\ll\lambda$ can be justified and Eqn. (10) becomes applicable. Using the same formulation, significant damping can then occur for even small values of $\eta_{i}$. As opposed to the first scenario where the value of $\eta$ dominates damping, this effect is present in geometries where the $(\lambda/r_{i})$ ratio is large, which amplifies the value of the negative exponent in Eqn. (10).
As an example, suppose that $r_{i}=3\times 10^{-28}$ m (a pGW source the size of the horizon scale) with $\lambda=3\times 10^{-22}$ m (the lower wavelength limit), then ${H}_{o}=0.5{H}_{i}$ for $\eta=2.0\times 10^{25}$ kg/m/s. This value is much smaller than that of Eqn. (16), and is of the same order as values considered for CCSNe. The timescale of the damping is $10^{-36}$ s, which is much shorter than the timescale of inflation.
@@ 4 Conclusion
This essay delves into the astrophysical contexts where GW damping and GW heating are highly relevant. Specifically, we examine the effects of a matter shell enveloping the GW source. Initially, we analyze a low-density dust shell, providing the modified GW strain equation (Eqn. 7). Expanding our discourse, we explore the implications of a viscous shell encircling the GW source. Energy dissipation occurs within this shell as GWs traverse it, leading to GW damping and heating of the shell.